# L-MTP: Leap Multi-Token Prediction Beyond Adjacent Context for Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have achieved notable progress. Despite their success, next-token prediction (NTP), the dominant method for LLM training and inference, is constrained in both contextual coverage and inference efficiency due to its inherently sequential process. To overcome these challenges, we propose leap multi-token prediction (L-MTP), an innovative token prediction method that extends the capabilities of multi-token prediction (MTP) by introducing a leap-based mechanism. Unlike conventional MTP, which generates multiple tokens at adjacent positions, L-MTP strategically skips over intermediate tokens, predicting non-sequential ones in a single forward pass. This structured leap not only enhances the model’s ability to capture long-range dependencies but also enables a decoding strategy specially optimized for non-sequential leap token generation, effectively accelerating inference. We theoretically demonstrate the benefit of L-MTP in improving inference efficiency. Experiments across diverse benchmarks validate its merit in boosting both LLM performance and inference speed. The source code will be publicly available.
@@ 1 Introduction
Large language models (LLMs) have demonstrated rapid and remarkable progress, driven by data, computing, and architectural innovation advances [1, 2, 3, 4, 5, 6, 7]. They exhibit strong capabilities in world knowledge acquisition [8, 9] and enable breakthroughs across a wide range of research domains, such as chemistry [10, 11], biology [12, 13], and medicine [14, 15]. As model scales and training data continue to increase, LLMs are attaining ever more powerful generalization and reasoning abilities [16, 17, 18, 19].
Next-token prediction (NTP) remains the mainstream strategy for both training and inference in LLMs [20, 21, 22, 23, 24, 25]. It generates tokens in an autoregressive manner, where each token is predicted based only on the preceding context (see Figure 2(1(a))). However, despite its conceptual simplicity, NTP results in inefficient generation, and limits the model to a focused yet short contextual horizon, and overlooks “hard” decisions [26, 27]. Intriguingly, LLMs are verified with inherent pre-planning capabilities, which indicates the potential of extending NTP to predict multiple tokens at once [27]. This gives rise to the multi-token prediction (MTP) paradigm [27] (see Figure 2(1(b))). Specifically, by incorporating additional language model heads, MTP enables the parallel prediction of a sequence of adjacent tokens and brings two key benefits. First, it provides a broader training signal by supervising multiple upcoming tokens at each step, which can enhance performance in tasks requiring long-range reasoning or planning [27, 28, 29]. Second, it enables faster inference by generating multiple tokens in a single forward pass, reducing latency and increasing throughput in applications with efficiency constraints [27].
@@
The study of L-MTP can be justified from both human thinking and recent trends in language model reasoning. In human thinking, we rarely reason in a strictly sequential fashion. Instead, we often skip over intermediate elements to complete reasoning more efficiently [30, 31, 32]. This leap-wise reasoning aligns naturally with L-MTP’s mechanism of skipping intermediate tokens and predicting non-adjacent ones. Similarly, in language model reasoning, recent advances in efficient reasoning have revealed that many intermediate reasoning steps can be compressed or abstracted without loss of correctness [33, 34, 35]. By predicting tokens at leaping positions, L-MTP mimics this abstraction process, not by explicitly modeling token importance, but by altering the prediction pattern to skip intermediate positions, to accelerate LLM inference.
@@ 2 Preliminaries
MTP. A natural extension for NTP is MTP [27] that predicts multiple tokens at once. Given input tokens $\mathrm{x}_{\leq t}$, the LLM with parameters $\bar{\theta}$, predicts the following $n$ tokens, by involving more output heads (e.g., 4 output heads totally). Therefore, the optimization objective is derived from Eq. (1) to:
@@
In this case, the LLM is requested to pre-plan the adjacent context rather than the single next token. For MTP, during decoding, the next $n$ tokens can be sampled independently, following $p(\mathrm{x}_{t+i}|\mathrm{x}_{\leq t};\bar{\theta}),i\in\{1,2,\dots,n\}$. Recent research [27, 36] disentangles the LLM with the LLM backbone and output heads, where the former yields hidden states and the latter maps them into a vocabulary distribution. Therefore, the LLM backbone can be equipped with multiple heads to predict the tokens independently according to the shared hidden states. To this end, we have
@@
Figure 2: MTP with self-speculative decoding.
where $\mathbf{z}$ denotes the hidden states, $\theta^{\prime}$ represents the parameters for the LLM backbone, and $\theta^{i}$ is the parameters for the $i$-th output head. Following token predictions, MTP typically incorporates verification and acceptance sub-procedures to determine which tokens are eligible for output (cf., Figure 2). Specifically, the verification sub-procedure invokes LLMs to evaluate the predictions in parallel and records their probabilities for determining the acceptance. Accepted token will be used to update the KV cache, and the process proceeds to the next iteration. Such decoding is proven as a lossless LLM acceleration technique, where the sampling is consistent with the distribution of vanilla autoregressive decoding [37].
@@ 3 L-MTP: Leap Multi-Token Prediction
Overview. Despite the potential of MTP, we go beyond it with one innovative solution to achieve a broader prediction range and faster inference. Unlike conventional MTP, which focuses on consecutive tokens, L-MTP introduces a leap-based strategy, allowing it to predict tokens at non-sequential positions within the context window. This design enables the model to efficiently capture long-range dependencies without the need for dense token predictions. During inference, L-MTP reuses partially overlapping context across prediction steps, maximizing information utilization while minimizing redundant calculations. Given input tokens $\mathrm{x}_{\leq t}$, L-MTP aims to predict a sequence of tokens at leaping intervals, specifically at positions, i.e., $\mathrm{x}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[t+k(n-1)+1,\dots,t+k+1,t+1]}}$, where $k$ denotes the number of jumped tokens. For example, with $t$ input tokens and $k=2$, the model is expected to predict the tokens at positions $[t+1,t+3,t+5,\dots,t+2n-1]$, effectively skipping intermediate tokens in each prediction step. We detail our L-MTP below.
@@ 3.2 L-MTP Inference Procedure
Looking backward. Given $\mathrm{x}_{\leq t}$ tokens [^3] Typically, $t>1$ with the input containing at least one start token, like “<|begin_of_text|>” in Llama series., L-MTP predicts tokens $\{\mathrm{x}_{t+k(i-1)+1}\}_{i\in[n]}$, while leaving gaps between them (i.e., $k-1$ tokens are skipped). However, if we look backward, the desired tokens have already been predicted by prior steps. For instance, tokens $\{\mathrm{x}_{t+k(i-1)}\}_{i\in[n]}$ are predicted given $\mathrm{x}_{\leq t-1}$. In this case, we have:
@@
Combining with tree attention. L-MTP seamlessly integrates with speculative decoding by sampling consecutive token sequences for verification. Drawing inspiration from parallel decoding [38, 39, 36, 40], we combine L-MTP with tree attention to enable efficient decoding. We construct a hierarchical tree structure, where the $i$-th layer represents candidate tokens generated by the $i$-th prediction head. Paths in the tree are explored to identify the accepted one. To facilitate parallel verification, we design a tree attention mask that restricts each hidden state to attend only to its ancestors [38, 36]. Figure 4 illustrates the implementation of the tree attention with L-MTP decoding. Further details are provided in Appendix C.2.
@@ 4 Theoretical Analyses
Remark. Attenuation reflects the increasing uncertainty in the language model’s predictions as it forecasts tokens further into the future. This behavior arises because the prediction of $\mathrm{x}_{t+i}$ relies on the fixed context $\mathrm{x}_{\leq t}$, and the influence of this context diminishes with increasing $i$. As a result, the model’s confidence, as measured by the marginal probability $p(\mathrm{x}_{t+i}|\mathrm{x}_{\leq t})$, decreases monotonically.
@@
where vanilla strategy predicts tokens $\mathrm{x}_{t+1},\mathrm{x}_{t+2},\dots,\mathrm{x}_{t+n}$ sequentially using the hidden state at $t$. L-MTP predicts two interleaved sequences. Specifically, L-MTP uses the hidden state at $t-1$ to compensate for the non-predicted tokens.
@@
Remark. L-MTP introduces a longer prediction range ($k(n-1)+1$) to compensate for the loss of confidence on leaping positions. Theorem 3 reveals the relation between attenuation and the number of prediction heads. Less attenuation indicates higher speed-up of L-MTP compared to the vanilla strategy. In practice, $n$ would not be too large. Less attenuation ($\gamma$) with fewer overheads of heads ($n$) leads to further higher speed up. We illustrate the simulated curves for a better understanding of the superiority of L-MTP (see Figure 5).
@@ 5.1 Experimental Setup
Datasets. We curate the training dataset from Math [41], Evol-Instruct-Code [42, 43], and Alpaca-GPT4 [44]. In the first stage, we use the full data for self-distillation. For the second stage, we randomly select 10,000 examples with a ratio of 4:4:2, corresponding to math, code, and general data, respectively. To benchmark the methods, we select Math500 [45] (4-shot) and GSM8K [46] (4-shot) for math evaluation, MBPP, MBPP+ [47, 48], HumanEval, and HumanEval+ [49, 48] for code evaluation, and MMLU [50] and IFEval [51] for general evaluation. We detail the statistics and utilization of these datasets in Appendix B.3, Appendix B.4, and Appendix B.6.
@@
Implementation details. To adapt L-MTP for NTP-based methods, we employ a two-stage training procedure. At the head warming up stage, we freeze the LLM backbone while training the heads with a learning rate of 1×$\times 10^{-3}$ for 5 epochs. We utilize the cosine scheduler and set the warmup ratio as 0.1. At the next stage, we utilize LoRA [52] with rank being 32 and alpha being 16 to tune the full model. Here we only train the model for 3 epochs with the learning rate being 1×$\times 10^{-5}$. We set $k=2$ and $n=4$ by default. This training setting is also employed for MTP implementation to ensure fairness. We also provide the pseudo-code of L-MTP in Appendix B.1. All the experiments are conducted on 2$\times$ NVIDIA H100-80G GPUs.
@@ 5.2 Results and Discussions
Viccuna 7B MTP 1.83$\times$ 1.97$\times$
L-MTP 2.32$\times$ 2.01$\times$
Viccuna 13B MTP 2.24$\times$ 1.98$\times$
L-MTP 2.43$\times$ 2.02$\times$
We also explore the potential by extending L-MTP decoding to existing models, like Medusa [36], which is specialized for improving the acceptance rate for multiple heads. We equip these models with L-MTP decoding and showcase the results in Table 2. Directly changing the decoding strategy to the leaping paradigm brings up to 1.3$\times$ speed up (22% relative boosting). These results demonstrate the potential of L-MTP, especially for models with higher acceptance rates.
@@
Potential analysis (RQ4). We emphasize the potential of L-MTP by investigating the myopia of LLMs and the effect of data amount. (1) Myopic generation. We demonstrate the prediction accuracy across different scales of models, as shown in Figure 9. The accuracy drops consistently when changing the small model to the larger one for all series of LLMs, indicating the inherent myopia imposed by NTP pre-training. We also provide the loss curves during training in Appendix D, which show an inflection when warming up the heads. Recent work [27] suggests training a model from scratch with the MTP objective. This is also promising for L-MTP to inherently benefit the model with a broader range of predictions and faster inference. (2) Data scales. We illustrate the increasing prediction accuracy when adding more data in Figure 9 at the head warming up stage. Large-scale data introduces more diversity to help additional heads adapt the LLM backbone. However, we also observe that the increase is not linear. To obtain higher accuracy, equipping L-MTP with more sophisticated techniques for training or model architecture [53, 40] will be also promising.
@@ 6.1 Multi-Token Prediction
Previous studies demonstrate that multi-token prediction (MTP) encourages pre-planning capabilities in large language models (LLMs). Qi et al. [26] pioneer $n$-step-ahead prediction to optimize language models, mitigating overfitting to strong local dependencies. Gloeckle et al. [27] pretrain LLMs with additional prediction heads to achieve significant performance improvements, particularly on code-related tasks. Industrial deployments have also adopted MTP to improve both training efficiency and pre-planning [28, 29]. MTP has sparked growing interest in exploring its potential, including adapting next-token prediction (NTP) models for MTP [54] and applying it to domains such as speech [55]. Furthermore, recent work investigates MTP’s potential for inference acceleration by incorporating additional prediction heads, as exemplified by Medusa [36]. We discuss LLM inference acceleration further in the next subsection.
In addition, recent research on MTP shows significant promise, with the support of LLMs inherently maintaining certain pre-planning [56]. These methods typically assume prediction within an adjacent context for LLMs, predicting the next $n$ tokens simultaneously at each time step. We go beyond its prediction pattern and introduce leaps between prediction tokens, therefore extending a broader training signal.

### v2

@@ Abstract
Large language models (LLMs) have achieved notable progress. Despite their success, next-token prediction (NTP), the dominant method for LLM training and inference, is constrained in both contextual coverage and inference efficiency due to its inherently sequential process. To overcome these challenges, we propose leap multi-token prediction (L-MTP), an innovative token prediction method that extends the capabilities of multi-token prediction (MTP) by introducing a leap-based mechanism. Unlike conventional MTP, which generates multiple tokens at adjacent positions, L-MTP strategically skips over intermediate tokens, predicting non-sequential ones in a single forward pass. This structured leap not only enhances the model’s ability to capture long-range dependencies but also enables a decoding strategy specially optimized for non-sequential leap token generation, effectively accelerating inference. We theoretically demonstrate the benefit of L-MTP in improving inference efficiency. Experiments across diverse benchmarks validate its merit in boosting both LLM performance and inference speed. The source code is available at https://github.com/Xiaohao-Liu/L-MTP.
@@ 1 Introduction
Large language models (LLMs) have demonstrated rapid and remarkable progress, driven by data, computing, and architectural innovation advances floridi2020gpt ; grattafiori2024llama ; team2025kimi ; chen2024expanding ; wang2025baichuan ; team2023gemini ; luo2024mmevol . They exhibit strong capabilities in world knowledge acquisition ge2024worldgpt ; yildirim2024task and enable breakthroughs across a wide range of research domains, such as chemistry boiko2023autonomous ; jablonka2024leveraging , biology lam2024large ; gao2024empowering , medicine thirunavukarasu2023large ; clusmann2023future , and personalization zhao2025teaching ; qiu2025latent ; liu2025fine ; qiu2025measuring . As model scales and training data continue to increase, LLMs are attaining ever more powerful generalization and reasoning abilities kaplan2020scaling ; ren2025deepseekproverv2 ; bi2024deepseek ; zeng2023glm ; zhao2025exploring ; chen2025emergence .
Next-token prediction (NTP) remains the mainstream strategy for both training and inference in LLMs yang2024qwen25 ; chen2024next ; zhang2025hierarchical ; he2024law ; flemings2024differentially ; nguyen2025turning . It generates tokens in an autoregressive manner, where each token is predicted based only on the preceding context (see Figure 2(1(a))). However, despite its conceptual simplicity, NTP results in inefficient generation, and limits the model to a focused yet short contextual horizon, and overlooks “hard” decisions qi2020prophetnet ; gloeckle2024better . Intriguingly, LLMs are verified with inherent pre-planning capabilities, which indicates the potential of extending NTP to predict multiple tokens at once gloeckle2024better . This gives rise to the multi-token prediction (MTP) paradigm gloeckle2024better  (see Figure 2(1(b))). Specifically, by incorporating additional language model heads, MTP enables the parallel prediction of a sequence of adjacent tokens and brings two key benefits. First, it provides a broader training signal by supervising multiple upcoming tokens at each step, which can enhance performance in tasks requiring long-range reasoning or planning gloeckle2024better ; liu2024deepseek ; xiaomi2025mimo . Second, it enables faster inference by generating multiple tokens in a single forward pass, reducing latency and increasing throughput in applications with efficiency constraints gloeckle2024better .
@@
The study of L-MTP can be justified from both human thinking and recent trends in language model reasoning. In human thinking, we rarely reason in a strictly sequential fashion. Instead, we often skip over intermediate elements to complete reasoning more efficiently reyna1995fuzzy ; brainerd2002fuzzy ; fiske1991social . This leap-wise reasoning aligns naturally with L-MTP’s mechanism of skipping intermediate tokens and predicting non-adjacent ones. Similarly, in language model reasoning, recent advances in efficient reasoning have revealed that many intermediate reasoning steps can be compressed or abstracted without loss of correctness lin2024rho ; chen2024not ; xia2025tokenskip . By predicting tokens at leaping positions, L-MTP mimics this abstraction process, not by explicitly modeling token importance, but by altering the prediction pattern to skip intermediate positions, to accelerate LLM inference.
@@ 2 Preliminaries
MTP. A natural extension for NTP is MTP gloeckle2024better that predicts multiple tokens at once. Given input tokens $\mathrm{x}_{\leq t}$, the LLM with parameters $\bar{\theta}$, predicts the following $n$ tokens, by involving more output heads (e.g., 4 output heads totally). Therefore, the optimization objective is derived from Eq. (1) to:
@@
In this case, the LLM is requested to pre-plan the adjacent context rather than the single next token. For MTP, during decoding, the next $n$ tokens can be sampled independently, following $p(\mathrm{x}_{t+i}|\mathrm{x}_{\leq t};\bar{\theta}),i\in\{1,2,\dots,n\}$. Recent research gloeckle2024better ; cai2024medusa disentangles the LLM with the LLM backbone and output heads, where the former yields hidden states and the latter maps them into a vocabulary distribution. Therefore, the LLM backbone can be equipped with multiple heads to predict the tokens independently according to the shared hidden states. To this end, we have
@@
Figure 2: MTP with self-speculative decoding, involving three sub-procedures.
where $\mathbf{z}$ denotes the hidden states, $\theta^{\prime}$ represents the parameters for the LLM backbone, and $\theta^{i}$ is the parameters for the $i$-th output head. Following token predictions, MTP typically incorporates verification and acceptance sub-procedures to determine which tokens are eligible for output (cf., Figure 2). Specifically, the verification sub-procedure invokes LLMs to evaluate the predictions in parallel and records their probabilities for determining the acceptance. Accepted token will be used to update the KV cache, and the process proceeds to the next iteration. Such decoding is proven as a lossless LLM acceleration technique, where the sampling is consistent with the distribution of vanilla autoregressive decoding leviathan2023fast .
@@ 3 L-MTP: Leap Multi-Token Prediction
Overview. Despite the potential of MTP, we go beyond it with one innovative solution to achieve a broader prediction range and faster inference. Unlike conventional MTP, which focuses on consecutive tokens, L-MTP introduces a leap-based strategy, allowing it to predict tokens at non-sequential positions within the context window. This design enables the model to efficiently capture long-range dependencies without the need for dense token predictions. During inference, L-MTP reuses partially overlapping context across prediction steps, maximizing information utilization while minimizing redundant calculations. Given input tokens $\mathrm{x}_{\leq t}$, L-MTP aims to predict a sequence of tokens at leaping intervals, specifically at positions, i.e., $\mathrm{x}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}[t+k(n-1)+1,\dots,t+k+1,t+1]}}$, where $k$ denotes the number of jumped tokens. For example, with $t$ input tokens and $k=2$, the model is expected to predict the tokens at positions $t+1,t+3,t+5,\dots,t+2n-1]$, effectively skipping intermediate tokens in each prediction step. We detail our L-MTP below.
@@ 3.2 L-MTP Inference Procedure
Looking backward. Given $\mathrm{x}_{\leq t}$ tokens [^3] Typically, $t>1$ with the input containing at least one start token, like “<|begin_of_text|>” in Llama series., L-MTP predicts tokens $\{\mathrm{x}_{t+k(i-1)+1}\}_{i\in[n]}$, while leaving gaps between them (i.e., k1k-1 tokens are skipped). However, if we look backward, the desired tokens have already been predicted by prior steps. For instance, tokens $\{\mathrm{x}_{t+k(i-1)}\}_{i\in[n]}$ are predicted given $\mathrm{x}_{\leq t-1}$. In this case, we have:
@@
Combining with tree attention. L-MTP seamlessly integrates with speculative decoding by sampling consecutive token sequences for verification. Drawing inspiration from parallel decoding miao2024specinfer ; sun2023spectr ; cai2024medusa ; li2024eagle , we combine L-MTP with tree attention to enable efficient decoding. We construct a hierarchical tree structure, where the $i$-th layer represents candidate tokens generated by the $i$-th prediction head. Paths in the tree are explored to identify the accepted one. To facilitate parallel verification, we design a tree attention mask that restricts each hidden state to attend only to its ancestors miao2024specinfer ; cai2024medusa . Figure 4 illustrates the implementation of the tree attention with L-MTP decoding. Further details are provided in Appendix C.2.
@@ 4 Theoretical Analyses
Remark. Attenuation reflects the increasing uncertainty in the language model’s predictions as it forecasts tokens further into the future. This behavior arises because the prediction of $\mathrm{x}_{t+i}$ relies on the fixed context $\mathrm{x}_{\leq t}$, and the influence of this context diminishes with increasing ii. As a result, the model’s confidence, as measured by the marginal probability $p(\mathrm{x}_{t+i}|\mathrm{x}_{\leq t})$, decreases monotonically.
@@
where vanilla strategy predicts tokens $\mathrm{x}_{t+1},\mathrm{x}_{t+2},\dots,\mathrm{x}_{t+n}$ sequentially using the hidden state at $t$. L-MTP predicts two interleaved sequences. Specifically, L-MTP uses the hidden state at t1t-1 to compensate for the non-predicted tokens.
@@
Remark. L-MTP introduces a longer prediction range (k(n1)+1k(n-1)+1) to compensate for the loss of confidence on leaping positions. Theorem 3 reveals the relation between attenuation and the number of prediction heads. Less attenuation indicates higher speed-up of L-MTP compared to the vanilla strategy. In practice, $n$ would not be too large. Less attenuation ($\gamma$) with fewer overheads of heads ($n$) leads to further higher speed up. We illustrate the simulated curves for a better understanding of the superiority of L-MTP (see Figure 5).
@@ 5.1 Experimental Setup
Datasets. We curate the training dataset from Math hendrycksmath2021 , Evol-Instruct-Code luowizardcoder ; codealpaca , and Alpaca-GPT4 peng2023instruction . In the first stage, we use the full data for self-distillation. For the second stage, we randomly select 10,000 examples with a ratio of 4:4:2, corresponding to math, code, and general data, respectively. To benchmark the methods, we select Math500 lightman2023let (4-shot) and GSM8K cobbe2021training (4-shot) for math evaluation, MBPP, MBPP+ austin2021program ; liu2023your , HumanEval, and HumanEval+ chen2021evaluating ; liu2023your for code evaluation, and MMLU hendrycks2020measuring and IFEval zhou2023instruction for general evaluation. We detail the statistics and utilization of these datasets in Appendix B.3, Appendix B.4, and Appendix B.6.
@@
[…]
@@ 5.2 Results and Discussions
Viccuna 7B
[…]
Viccuna 13B
[…]
@@
Potential analysis (RQ4). We emphasize the potential of L-MTP by investigating the myopia of LLMs and the effect of data amount. (1) Myopic generation. We demonstrate the prediction accuracy across different scales of models, as shown in Figure 9. The accuracy drops consistently when changing the small model to the larger one for all series of LLMs, indicating the inherent myopia imposed by NTP pre-training. We also provide the loss curves during training in Appendix D, which show an inflection when warming up the heads. Recent work gloeckle2024better suggests training a model from scratch with the MTP objective. This is also promising for L-MTP to inherently benefit the model with a broader range of predictions and faster inference. (2) Data scales. We illustrate the increasing prediction accuracy when adding more data in Figure 9 at the head warming up stage. Large-scale data introduces more diversity to help additional heads adapt the LLM backbone. However, we also observe that the increase is not linear. To obtain higher accuracy, equipping L-MTP with more sophisticated techniques for training or model architecture anknerhydra ; li2024eagle will also be promising.
@@ 6.1 Multi-Token Prediction
Previous studies demonstrate that multi-token prediction (MTP) encourages pre-planning capabilities in large language models (LLMs). Qi et al. qi2020prophetnet pioneer $n$-step-ahead prediction to optimize language models, mitigating overfitting to strong local dependencies. Gloeckle et al. gloeckle2024better pretrain LLMs with additional prediction heads to achieve significant performance improvements, particularly on code-related tasks. Industrial deployments have also adopted MTP to improve both training efficiency and pre-planning liu2024deepseek ; xiaomi2025mimo . MTP has sparked growing interest in exploring its potential, including adapting next-token prediction (NTP) models for MTP mehra2025multi and applying it to domains such as speech wang2025vocalnet . Furthermore, recent work investigates MTP’s potential for inference acceleration by incorporating additional prediction heads, as exemplified by Medusa cai2024medusa . We discuss LLM inference acceleration further in the next subsection.
In addition, recent research on MTP shows significant promise, with the support of LLMs inherently maintaining certain pre-planning wulanguage . These methods typically assume prediction within an adjacent context for LLMs, predicting the next $n$ tokens simultaneously at each time step. We go beyond its prediction pattern and introduce leaps between prediction tokens, therefore extending a broader training signal.